day shift, 28Sept99

Marj Corcoran (corcoran@physics.rice.edu)
Tue, 28 Sep 1999 15:55:42 -0500 (CDT)


Shift Summary for day shift, 28 Sept 99

On Shift: M. Corcoran and Rick Kessler

Hours of beam available at all intensities: 5.2
Hours of beam available at nominal intensity: 5.2
Hours of E799 data (special runs not included): 2.7
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Swap HV divider on 1533; reboot atev1; multiple
drt_fresh_starts, restart sdaq processes

========================================================================

Runs ended this shift:
Run 14753 E799 data
Run 14754 E799 data in progress

=======================================================================

Tapes brought to FCC this shift: UPA,B,C 40-42 and UPS14

==================================================================

Description of shift: Start shift with divider swap, then much trouble
starting the run. KFREND was way slow, reboot atev1, restart sdaq processes,
but they didn't all start. Had to page Theo. Drt_fresh_start twice.
Run is finally started around 9:30. Smooth running after that, when there was
beam.

Beam off for 1.5 hrs due to power supply problem, and another 1.5 hrs
due to a thunderstorm. Occasional aborts.

===================================================================

run spill time comment
-----------------------------------------------------------------
14753 1-22 9:30 sdaq was not getting sem or spill number; page Theo
problem is fixed around spill 22

14753 15 9:35 BobH notices beam is off horizontally, call MCR

14753 20 9:40 Beam OK now...autotune restarted

145753 30 9:55 TeV power supply problem, no est, beam gone

14753 90 11:20 Beam returns

12:25 ONSPLIT requests tape UPX003 on ktev3, tps2d3,
RK complies

14753 164 12:56 MCR calls, lightening has killed interlocks on
neutrino primary. No estimate.

13:00 ONSPLIT requests UPH004 on ktev4, tps2d3, we comply

14753 178 13:00 Estimate for beam 1 hr

14753 228 14:20 End run, as there is still no beam

14754 1 14:25 Ending the previous run did the trick, beam is back,
start new run!

15:40 End of shift, but no new shiftees!

15:50 Still no evening shift...page Sydney who says he
is actually not on shift. Rick F who happens to be
around graciously agrees to stay until someone
arrives.

15:55 Val arrives! I declare the day shift over

================================================================

Interesting factoid: 1kg of antimatter (with 1 kg of matter) could
supply the US consumption of elecricity for a week! 8 tons of uranium
will do the same...or 30 million tons of coal. For more fascinating facts,
attend Rick K's lecture at the Adler Planatarium on Oct 13, 6:30 PM!

===============================================================

Follow up on next shift: Channel 961 has comparator problems at a low
rate...24 at spill 134 in run 14753. May want to take some action on down
day.